Emmanuel and Camille are married, file jointly, and have two children, ages three and seven, whom they support. Their AGI is $28,470.

Click here to access Exhibit 13.3 Earned Income Credit and Phaseout Percentages.

Carry out computations to two decimal places, and round your final answer to the nearest dollar.

Enter the amount of Emmanuel and Camille's earned income credit: $_______Exhibit 13.3 Child and Dependent Care Credit Computations Adjusted Gross Income Applicable Over But Not Over $15,000 0 15,000
Answer #1
Assumed year: 2019
Maximum Earned Income Credit allowed to be taken as per Rules                                 5,828.00
Credit phase-out begins at $24,820 and the tax credit reduction rate is 21.06%
Now, AGI reported (a)                              28,470.00
Credit phase-out AGI (b)                              24,820.00
Net amount on which the reduction will apply = c = (a-b)                                 3,650.00
Rate of credit phase-out (d) 21.06%
Amount of credit phased-out = e = c X d                                    768.69
Net Amount of Earned Income credit allowed = Maximum credit allowed to be taken as per Rules - Amount of credit phased out = 5828 - 768.69 = 5059.31 = $5,059